OpenStudy (anonymous):

derivative of y=e^(5x)-sin^2(8x)

OpenStudy (anonymous):

chain rule.

OpenStudy (anonymous):

yes but how do u go about doing it? like which one is the outside eq?

OpenStudy (anonymous):

Okay first of all, we need to do each term separately.

OpenStudy (anonymous):

\[ e^{5x} \]Let \(u = 5x\). Now we can say: \[ d(e^{5x}) =d(e^u) = e^u\;du \]Do you follow so far?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[ e^udu = e^{5x}d(5x) = e^{5x}5dx \]

OpenStudy (anonymous):

Put it all together, and we get: \[ d(e^{5x}) = 5e^{5x}dx \implies \frac{d(e^{5x})}{dx} = 5e^{5x} \]

OpenStudy (anonymous):

uhh what

OpenStudy (anonymous):

ohh kk

OpenStudy (anonymous):

Now, do you think you can do the same process for \[ -\sin^2(8x) \]

OpenStudy (anonymous):

In this case, I suggest \(u = \sin(8x)\). And if you have trouble with \(du\), then let \(v=8x\).

OpenStudy (anonymous):

so cosv-sinv?

OpenStudy (anonymous):

Okay, first step: \[ d(-\sin^2(8x)) = \ldots \]What is the first substitution?

OpenStudy (anonymous):

What should we choose for \(u=f(x)\)?

OpenStudy (anonymous):

uhh im getting more confused

OpenStudy (anonymous):

8x?

OpenStudy (anonymous):

Okay... Then \[ d(-\sin^2(8x)) = d(-\sin^2(u)) \]We have mad things a little easier. What else can we do to make it easier?

OpenStudy (anonymous):

change -sin^2 in two parts?

OpenStudy (anonymous):

Here is a hint: \[ d(-\sin^2(u)) = d(-[\sin(u)]^2) \]

OpenStudy (anonymous):

-2sin(u)

OpenStudy (anonymous):

ok, so \[ d(-\sin^2(u)) = -2\sin(u)d(\sin(u)) = -2\sin(u)\cos(u)du \]

OpenStudy (anonymous):

To make it more clear what I did: \[ d(-\sin^2(u)) = d(-v^2) = -2vdv = -2\sin(u)d(\sin(u)) = -2\sin(u)\cos(u)du \]

OpenStudy (anonymous):

It is a bit trickier when you do it without actually substituting the variable, but it still works!

OpenStudy (anonymous):

so the final without subtituting would be 5e^5x-2sin(8x)*cos(sin(8x))?

OpenStudy (anonymous):

Hold on...

OpenStudy (anonymous):

This is a simple double angle formula:\[ -2\sin(u)\cos(u)du = -\sin(2u)du \]

OpenStudy (anonymous):

Next we have: \[ -\sin(2u)du = -\sin(2(8x))d(8x) = -\sin(16x)8dx \]

OpenStudy (anonymous):

cause when i attempted, it said 5e^(5x)-16\cos(8x)*\sin(8x)\] was not the answer

OpenStudy (anonymous):

The answer is \[ 5e^{5x}-8\sin(16x) \]And if you didn't do the double angle simplification, it would be \[ 5e^{5x}-16\sin(8x)\sin(8x) \]

OpenStudy (anonymous):

neither of those were correct

OpenStudy (anonymous):

that was the correct answer^

OpenStudy (anonymous):

What was the original problem?

OpenStudy (anonymous):

The problem... is the minus sign.

OpenStudy (anonymous):

It was multiplication, not subtraction... that is why we got the wrong answer.